# Adaptive Informed Deep Neural Networks for Power Flow Analysis

### v1

@@ Abstract
This study introduces PINN4PF, an end-to-end deep learning architecture for power flow (PF) analysis that effectively captures the nonlinear dynamics of large-scale modern power systems. The proposed neural network (NN) architecture consists of two important advancements in the training pipeline: (A) a double-head feed-forward NN that aligns with PF analysis, including an activation function that adjusts to active and reactive power consumption patterns, and (B) a physics-based loss function that partially incorporates power system topology information. The effectiveness of the proposed architecture is illustrated through 4-bus, 15-bus, 290-bus, and 2224-bus test systems and is evaluated against two baselines: a linear regression model (LR) and a black-box NN (MLP). The comparison is based on (i) generalization ability, (ii) robustness, (iii) impact of training dataset size on generalization ability, (iv) accuracy in approximating derived PF quantities (specifically line current, line active power, and line reactive power), and (v) scalability. Results demonstrate that PINN4PF outperforms both baselines across all test systems by up to two orders of magnitude not only in terms of direct criteria, e.g., generalization ability but also in terms of approximating derived physical quantities.
@@
physics-informed neural networks, multilayer perceptron, voltage calculation, load flow, power systems.
@@ I Introduction
PF analysis can be performed by solving nonlinear and non-convex algebraic equations derived from the nodal balance of active and reactive power at each bus in power systems [3]. Exact analytical solutions for these equations, which also involve impedance parameters, load characteristics, and generator conditions of power systems, are impossible. Therefore, iterative numerical methods, such as the Gauss-Seidel and Newton-Raphson (NR) methods, are conventionally employed to converge to a solution that satisfies the PF equations within specified accuracy limits. Eventually, the solutions yield voltage phasors across the entire power system and, hence, provide a comprehensive understanding of the operational state of the power system [4, 5, 6].
@@
Conventional iterative numerical methods, however, face computational challenges in large-scale modern power systems. They struggle to capture the intricacies of power systems as a result of uncertainties involved, such as inaccurate line profiles due to weather conditions and aging, different types of loads, but also missing data on renewable energy resources [7]. Ineffective PF analysis under these circumstances can lead to safety threats, including renewable energy generation curtailment and blackouts, and difficulties in accommodating distributed energy resources [8, 9]. Addressing these challenges necessitates developing new approaches for PF analysis that are both computationally efficient and numerically stable.
@@
Deep learning approaches, and more specifically, artificial neural networks (NNs), are currently the most powerful set of numerical tools to provide accurate approximations of nonlinear problems [10]. Several studies have demonstrated the superiority of deep learning approaches in PF analysis in terms of computational time by orders of magnitude. At the same time, the accuracy of the solutions is competitive compared to the conventional iterative numerical methods (e.g., [11, 12]). NNs, therefore, can address the challenges mentioned above by leveraging the availability of massive measurements and/or augmented data, learning complex input-output relationships that are often difficult or even impossible for conventional iterative numerical methods to comprehend, and achieving the accuracy required for real-world applications [13]. Nevertheless, NNs are subject to overfitting, lack of generalization, and scalability issues. They are very unlikely to meet the physical constraints. Moreover, their performance relies heavily on the training dataset size and quality. In contrast, not enough data is always available due to privacy reasons and the presence of missing data, among others [14].
@@
Several studies in the literature have investigated the impact of various modifications on the efficacy of deep learning-based PF analysis. These modifications are categorized into three main stages: (i) pre-training (e.g., [15]); (ii) training (e.g., [16]); and (iii) post-training (e.g., [17]), as outlined in Table I. A majority of these studies have primarily focused on the training stage. The table also highlights instances where topology information, such as line physical properties (e.g., [18]) and graph topology (e.g., [19]) has been integrated into the training stage. The literature review also indicates the utilization of different prior knowledge for deep learning-based PF analysis and optimal power flow (OPF), including equality and inequality constraints (e.g., [20]) and PF equations (e.g., [21]). Among the three presented stages, this work focuses on the training stage.
@@
Pre-training Lei et al. (2021) [15] OPF $p_{d}$, $q_{d}$ $p_{g}$, $q_{g}$, $|v_{i}|$, $\delta_{i}$ Integrating measurement data into the pre-trained NN
@@
Hu et al. (2021) [22] PF $p_{d}$, $q_{d}$, $p_{g}$, $|v_{g}|$ $|v_{i}|$, $\delta_{i}$ Adding regularization terms based on structure of AC PF equations and topology of power system
@@
Lin et al. (2023) [23] PF $p_{d}$, $q_{d}$, $p_{g}$, $|v_{g}|$, $l_{ij}$, $\mathcal{G}$ $p_{i}$, $q_{i}$, $|v_{i}|$, $\delta_{i}$ Combining message passing graph NNs and high-order graph convolutional NNs
Liu et al. (2024) [24] PF $p_{d}$, $q_{d}$, $l_{ij}$ $v_{i}$ Using a physics-inspired structure integrating physical laws of PF
Nellikkath & Chatzivasileiadis (2021) [25] OPF $p_{d}$ $p_{g}$ Adding a penalty term of active power generation and consumption to loss function
Nellikkath & Chatzivasileiadis (2022) [26] OPF $p_{d}$, $q_{d}$ $p_{g}$, $q_{g}$ Adding a penalty term of active and reactive power generation and voltage magnitude to loss function
Hu & Zhang (2023) [27] OPF $p_{d}$, $q_{d}$ $|v_{i}|$, $\delta_{i}$ Developing an activation function and a modified loss function to satisfy physical constraints
@@
Post-training Donti et al. (2021) [17] OPF $p_{d}$, $q_{d}$, $|v_{r}|$ $p_{g}$, $|v_{r,g}|$ Implementing completion idea (hard constraints)
Pan et al. (2022) [29] OPF $p_{d}$, $q_{d}$ $p_{g}$, $q_{g}$, $|v_{i}|$, $\delta_{i}$ Implementing completion idea (hard constraints)
Li et al. (2023) [30] OPF $p_{d}$ $p_{g}$, $\delta_{i}$ Filtering generated values to ensure feasibility by comparing with actual values
$p_{i}$=nodal active power, $q_{i}$=nodal reactive power, $s_{ij}$=line apparent power, $|v|$=voltage magnitude, $\delta$=voltage angle, $\mathcal{G}$=graph topology, $l_{ij}$=line physical properties, $A$=Adjacency matrix, $p_{g}$=generator active power, $q_{g}$=generator reactive power, $p_{d}$=load active power, $q_{d}$=load reactive power.
Measurement data has been used in [15] to enhance deep learning-based OPF. While this modification has demonstrated improvements in accuracy, its applicability is confined to the pre-training stage. Physical properties have been used on a few occasions in the past to enhance deep learning-based PF analysis (e.g., [18, 31]), where the focus was mainly on the training stage. However, the high dependency on the physical properties of power systems affects the efficiency of this modification for deep learning-based PF analysis, certainly because physical properties are not always reliable due to different reasons, including aging and environmental conditions. In addition, integrating all the physical properties makes training processes computationally very expensive. The completion idea and filtering technique have also been employed in a few studies (e.g., [30]), where the focus is mainly on the post-training stage. Yet, similar to pre-training stage modifications, this approach has limitations as it is detached from the training stage. Among various methods, modified loss functions have been more widely used for deep learning-based PF analysis (e.g., [22]) and OPF (e.g., [27]). Physics-based loss functions have shown significant promise in enhancing the performance and reliability of NNs for PF and OPF.
This paper proposes an end-to-end deep learning architecture, hereafter called PINN4PF, to develop scalable and physics-informed NNs for PF analysis. To the best of our knowledge, such an architecture is currently not available in the scientific literature. This paper, therefore, focuses on the training stage and extends the existing knowledge of physics-informed NNs for PF analysis, e.g., [16, 22]. The main contributions are
@@
    developing a double-head feed-forward NN that aligns with PF analysis and includes an activation function that adjusts to to active and reactive power consumption patterns, i.e., input features, and
@@
This architecture can be a valuable addition to the literature and an important contribution as it provides a comprehensive understanding of different modifications to enhance the performance of NNs for PF analysis. Such understanding is required for researchers and practitioners to leverage the full potential of NNs and ensure reliable and efficient deep learning-based PF analysis. The application of PINN4PF is illustrated through four test power systems, namely 4-bus [32], 15-bus [33], 290-bus [34], and 2224-bus [35] test systems.
@@ II Power Flow Analysis
PF analysis aims to specify the state variables of power systems, i.e., $[\delta\text{ }v]^{T}$ where $\delta$ and $v$ denote the voltage phase angle and voltage magnitude, respectively. The AC PF analysis problem is a representation of Kirchhoff’s laws and is formulated in rectangular coordinates [22] as
$p_{i}$ $=\sum_{j=1}^{n}g_{ij}(\mu_{i}\mu_{j}+\omega_{i}\omega_{j})+b_{ij}(\omega_{i}\mu_{j}-\mu_{i}\omega_{j}),$ (1)
$q_{i}$ $=\sum_{j=1}^{n}g_{ij}(\omega_{i}\mu_{j}-\mu_{i}\omega_{j})-b_{ij}(\mu_{i}\mu_{j}+\omega_{i}\omega_{j}),$ (2)
where $i$ and $j$ are the indices of the buses, $n$ is the total number of buses in the power system, $p_{i}$ and $q_{i}$ are the active and reactive power injection/consumption at bus $i$, $g_{ij}$ and $b_{ij}$ are the real and imaginary components of the admittance $y_{ij}$ between buses $i$ and $j$, and $\mu_{i}=v_{i}cos\delta_{i}$ and $\omega_{i}=v_{i}sin\delta_{i}$ are the real and imaginary components of the voltage phasor at bus $i$. Here, $v_{i}$ and $\delta_{i}$ are the voltage magnitude and phase angle at bus $i$.
There are three types of buses in power systems: (i) reference bus, (ii) load bus (pq bus), and (iii) generation bus (pv bus). $v_{i}=1$ and $\delta_{i}=0$ are known, and $p_{i}$ and $q_{i}$ are unknown for the reference bus. For load buses, $p_{i}$ and $q_{i}$ are known, and $v_{i}$ and $\delta_{i}$ are unknown. Finally, $v_{i}$ and $p_{i}$ are known for generation buses, while $q_{i}$ and $\delta_{i}$ are unknown. This study considers cases with one reference bus and load buses for simplicity. For a power system consisting of a reference bus and load buses, a set of PF equations with the same number of equations and unknowns are achieved [36]
$p_{i}^{d}-p_{i}$ $=0,$ (3)
$q_{i}^{d}-q_{i}$ $=0,$ (4)
where $p_{i}$ and $q_{i}$ are defined by Equations (1)–(2), respectively, and $p_{i}^{d}$ and $q_{i}^{d}$ are active and reactive power injection/consumption at bus $i$, respectively.
Equations (3)–(4) are conventionally solved iteratively to specify the state of the power system $[\delta\text{ }v]^{T}$ until a convergence criterion is met, i.e., the mismatch between $p_{i}$ and $p_{i}^{d}$ but also $q_{i}$ and $q_{i}^{d}$ are small enough.
@@ III Deep Learning-based Power Flow Analysis
Deep learning-based PF analysis refers to developing NNs to approximate the variable states of power systems based on given historical system operation data, hereafter called dataset. The dataset includes input features $\vec{x}$ and output labels $\vec{y}$. For a power system with a reference bus and load buses, the input features are known variables, i.e., active and reactive power consumption at load buses $\vec{x}=\{(\vec{p}^{d}_{i},\vec{q}^{d}_{i}):i=1,2,\dots,n\}$. At the same time, the output labels are unknown variables, i.e., real and imaginary parts of voltages at load buses $\vec{y}=\{(\vec{\mu}_{i},\vec{\omega}_{i}):i=1,2,\dots,n\}$. The dimension of $\vec{x}$ and $\vec{y}$ is therefore $n\times 2$, where $n$ is the number of load buses. Note that the voltage magnitude and phase angle are known for the reference bus $i=0$, and the active and reactive power are unknown. Having voltages at all load buses approximated, the active and reactive power at the reference bus can be calculated.
The training of NNs is an iterative process and involves four steps. In the first iteration, the set of trainable parameters of the NN, i.e., the weight matrices and bias vectors $\theta=\{(W_{k},b_{k}):k=1,2,\dots,m\}$, are initialized in the update step, where $m$ is the number of hidden layers. In the forward step, the network is developed using Equations (6)–(8). The deviations of the approximated output obtained by the NN $\hat{\vec{y}}$ from the output labels $\vec{y}$ are computed in the loss step. Finally, the gradient of the deviations is calculated with respect to $\theta$ in the backward step. For the next iteration, $\theta$ is fine-tuned in the update step to reduce the deviations. The process continues until the maximum number of epochs is reached. In practice, the four steps can be individually and jointly modified to improve the overall performance of NNs. For example, it can be done by enhancing the network architecture in the forward step or adding a physical penalty term to the loss function in the loss step.

### v2

@@ Abstract
This study introduces PINN4PF, an end-to-end deep learning architecture for power flow (PF) analysis that effectively captures the nonlinear dynamics of large-scale modern power systems. The proposed neural network (NN) architecture consists of two important advancements in the training pipeline: (A) a double-head feed-forward NN that aligns with PF analysis, including an activation function that adjusts to the net active and reactive power injections patterns, and (B) a physics-based loss function that partially incorporates power system topology information. The effectiveness of the proposed architecture is illustrated through 4-bus, 15-bus, 290-bus, and 2224-bus test systems and is evaluated against two baselines: a linear regression model (LR) and a black-box NN (MLP). The comparison is based on (i) generalization ability, (ii) robustness, (iii) impact of training dataset size on generalization ability, (iv) accuracy in approximating derived PF quantities (specifically line current, line active power, and line reactive power), and (v) scalability. Results demonstrate that PINN4PF outperforms both baselines across all test systems by up to two orders of magnitude not only in terms of direct criteria, e.g., generalization ability, but also in terms of approximating derived physical quantities.
@@
physics-informed neural networks, multilayer perceptron, voltage calculation, load flow, state estimation.
@@ I Introduction
PF analysis can be performed by solving nonlinear and non-convex algebraic equations derived from the nodal balance of the net active and reactive power injections at each bus in power systems [3]. Exact analytical solutions for these equations, which also involve impedance parameters, load characteristics, and generator conditions of power systems, are impossible. Therefore, iterative numerical methods, such as the Gauss-Seidel and Newton-Raphson (NR) methods, are conventionally employed to converge to a solution that satisfies the PF equations within specified accuracy limits. Eventually, the solutions yield voltage phasors across the entire power system and, hence, provide a comprehensive understanding of the operational state of the power system [4, 5, 6].
@@
Conventional iterative numerical methods, however, face computational challenges in large-scale modern power systems. They struggle to capture the intricacies of power systems due to uncertainties involved, such as inaccurate line profiles due to weather conditions and aging, different types of loads, but also missing data on renewable energy resources [7]. Ineffective PF analysis under these circumstances can lead to safety threats, including renewable energy generation curtailment and blackouts, and difficulties in accommodating distributed energy resources [8, 9]. Addressing these challenges necessitates developing new approaches for PF analysis that are both computationally efficient and numerically stable.
@@
Deep learning approaches, and more specifically, artificial neural networks (NNs), are currently the most powerful set of numerical tools to provide accurate approximations of nonlinear problems (e.g., [10, 11, 12]). Several studies have demonstrated the superiority of deep learning approaches in PF analysis in terms of computational time by orders of magnitude (e.g., [13, 14, 15]). At the same time, the accuracy of the solutions is competitive compared to the conventional iterative numerical methods (e.g., [16, 17]). NNs, therefore, can address the challenges mentioned above by leveraging the availability of massive measurements and/or augmented data, learning complex input-output relationships that are often difficult or even impossible for conventional iterative numerical methods to comprehend, and achieving the accuracy required for real-world applications [18]. Nevertheless, NNs are subject to overfitting, the lack of generalization, and scalability issues. They are very unlikely to meet the physical constraints. Moreover, their performance relies heavily on the training dataset size and quality. In contrast, not enough data is always available due to privacy reasons and the presence of missing data, among others [19].
@@
Several studies in the literature have investigated the impact of various modifications on the efficacy of deep learning-based PF analysis (e.g., [20, 21]). These modifications are categorized into three main stages: (i) pre-training (e.g., [22]); (ii) training (e.g., [23]); and (iii) post-training (e.g., [24]), as outlined in Table I. A majority of these studies have primarily focused on the training stage. The table also highlights instances where topology information, such as line physical properties (e.g., [25]) and graph topology (e.g., [26]) has been integrated into the training stage. The literature review also indicates the utilization of different prior knowledge for deep learning-based PF analysis and optimal power flow (OPF), including equality and inequality constraints (e.g., [27]) and PF equations (e.g., [28]). Among the three presented stages, this work focuses on the training stage.
@@
Pre-training Lei et al. (2021) [22] OPF $p^{d}$, $q^{d}$ $p^{g}$, $q^{g}$, $|v_{i}|$, $\delta_{i}$ Integrating measurement data into the pre-trained NN
@@
Hu et al. (2021) [29] PF $p^{d}$, $q^{d}$, $p^{g}$, $|v_{g}|$ $|v_{i}|$, $\delta_{i}$ Adding regularization terms based on structure of AC PF equations and topology of power system
@@
Lin et al. (2023) [30] PF $p^{d}$, $q^{d}$, $p^{g}$, $|v_{g}|$, $l_{ij}$, $\mathcal{G}$ $p_{i}$, $q_{i}$, $|v_{i}|$, $\delta_{i}$ Combining message passing graph NNs and high-order graph convolutional NNs
Liu et al. (2024) [31] PF $p^{d}$, $q^{d}$, $l_{ij}$ $v_{i}$ Using a physics-inspired structure integrating physical laws of PF
Nellikkath & Chatzivasileiadis (2021) [32] OPF $p^{d}$ $p^{g}$ Adding a penalty term of active power generation and consumption to loss function
Nellikkath & Chatzivasileiadis (2022) [33] OPF $p^{d}$, $q^{d}$ $p^{g}$, $q^{g}$ Adding a penalty term of active and reactive power generation and voltage magnitude to loss function
Hu & Zhang (2023) [34] OPF $p^{d}$, $q^{d}$ $|v_{i}|$, $\delta_{i}$ Developing an activation function and a modified loss function to satisfy physical constraints
@@
Post-training Donti et al. (2021) [24] OPF $p^{d}$, $q^{d}$, $|v_{r}|$ $p^{g}$, $|v_{r,g}|$ Implementing completion idea (hard constraints)
Pan et al. (2022) [36] OPF $p^{d}$, $q^{d}$ $p^{g}$, $q^{g}$, $|v_{i}|$, $\delta_{i}$ Implementing completion idea (hard constraints)
Li et al. (2023) [37] OPF $p^{d}$ $p^{g}$, $\delta_{i}$ Filtering generated values to ensure feasibility by comparing with actual values
$p_{i}$=nodal active power, $q_{i}$=nodal reactive power, $s_{ij}$=line apparent power, $|v|$=voltage magnitude, $\delta$=voltage angle, $\mathcal{G}$=graph topology, $l_{ij}$=line physical properties, $A$=Adjacency matrix, $p^{g}$=generator active power, $q^{g}$=generator reactive power, $p^{d}$=load active power, $q^{d}$=load reactive power.
Measurement data has been used in [22] to enhance deep learning-based OPF. While this modification has demonstrated improvements in accuracy, its applicability is confined to the pre-training stage. Physical properties have been used on a few occasions in the past to enhance deep learning-based PF analysis (e.g., [25, 38]), where the focus was mainly on the training stage. However, the high dependency on the physical properties of power systems affects the efficiency of this modification for deep learning-based PF analysis, certainly because physical properties are not always reliable due to different reasons, including aging and environmental conditions. In addition, integrating all the physical properties makes training processes computationally very expensive. The completion idea and filtering technique have also been employed in a few studies (e.g., [37]), where the focus is mainly on the post-training stage. Yet, similar to pre-training stage modifications, this approach has limitations as it is detached from the training stage. Among various methods, modified loss functions have been widely used for deep learning-based PF analysis (e.g., [29]) and OPF (e.g., [34]). Physics-based loss functions have shown significant promise in enhancing the performance and reliability of NNs for PF and OPF.
This paper proposes an end-to-end deep learning architecture, hereafter called PINN4PF, to develop scalable and physics-informed NNs for PF analysis. The focus is on the training stage and extends the existing knowledge of physics-informed NNs for PF analysis, e.g., [23, 29]. The main contributions are:
@@
    developing a double-head feed-forward NN that aligns with PF analysis and includes an activation function that adjusts to the net active and reactive power injections patterns, i.e., input features, and
@@
The proposed PINN4PF architecture is designed to predict complex bus voltages in power systems by learning the mapping from the net active ($p$) and reactive ($q$) power injections to voltage states. It employs a double-head feed-forward NN structure, where one head predicts the real component ($\mu$) and the other predicts the imaginary component ($\omega$) of the complex voltages. Each head utilizes an adaptive activation function, a modified ReLU, whose slope coefficient is a trainable parameter. This design allows the activation functions to adapt dynamically during training via backpropagation, which ultimately enables the NN to learn complex relationships between the net power injections and voltage states effectively. The application of PINN4PF is illustrated through four test power systems, namely 4-bus [39], 15-bus [40], 290-bus [41], and 2224-bus [42] test systems.
@@ II Power Flow Analysis
Power flow (PF) analysis aims to specify the state variables of power systems, i.e., $[\delta\,v]^{T}$, where $\delta$ and $v$ denote the voltage phase angle and magnitude, respectively. The AC PF analysis problem is a representation of Kirchhoff’s laws and is formulated in rectangular coordinates [29] as:
$p_{i}$ $=\sum_{j=1}^{n}g_{ij}(\mu_{i}\mu_{j}+\omega_{i}\omega_{j})+b_{ij}(\omega_{i}\mu_{j}-\mu_{i}\omega_{j}),$ (1a)
$q_{i}$ $=\sum_{j=1}^{n}g_{ij}(\omega_{i}\mu_{j}-\mu_{i}\omega_{j})-b_{ij}(\mu_{i}\mu_{j}+\omega_{i}\omega_{j}),$ (1b)
where $i$ and $j$ are the indices of the buses, $n$ is the total number of buses in the power system, $p_{i}$ and $q_{i}$ are the net active and reactive power injections at bus $i$, $g_{ij}$ and $b_{ij}$ are the real and imaginary components of the admittance $y_{ij}$ between buses $i$ and $j$, $\mu_{i}=v_{i}cos\delta_{i}$ and $\omega_{i}=v_{i}sin\delta_{i}$ are the real and imaginary components of the voltage phasor at bus $i$. Here, $v_{i}$ and $\delta_{i}$ are the voltage magnitude and phase angle at bus $i$.
There are three types of buses in power systems: (i) reference bus, (ii) load bus (pq bus), and (iii) generation bus (pv bus). $v_{i}$ and $\delta_{i}$ are known, and $p_{i}$ and $q_{i}$ are unknown for the reference bus. For load buses, $p_{i}$ and $q_{i}$ are known, and $v_{i}$ and $\delta_{i}$ are unknown. $v_{i}$ and $p_{i}$ are known for generation buses, while $q_{i}$ and $\delta_{i}$ are unknown.
This study considers cases with one reference bus and load buses for simplicity. A standard approach to handling pv buses is converting pv buses to pq buses by estimating the reactive power while applying constraints to maintain voltage levels within acceptable limits. Once converted, pv buses can be treated as pq buses for training. It should also be noted that while a voltage magnitude of $1.0$ p.u. and a phase angle of 00 degrees are assumed in this study, it is flexible to accommodate different reference bus settings.
For a power system consisting of a reference bus and load buses, a set of PF equations with the same number of equations and unknowns is achieved [43]:
$p_{i}^{d}-p_{i}$ $=0,$ (2a)
$q_{i}^{d}-q_{i}$ $=0,$ (2b)
where $p_{i}$ and $q_{i}$ are defined by (1a) and (1b), respectively, and $p_{i}^{d}$ and $q_{i}^{d}$ are the net active and reactive power injections at bus $i$, respectively. (2) is conventionally solved iteratively to specify $[\delta\,v]^{T}$ until a convergence criterion is met, i.e., the mismatch between $p_{i}$ and $p_{i}^{d}$ but also $q_{i}$ and $q_{i}^{d}$ are small enough.
@@ III Deep Learning-based Power Flow Analysis
Deep learning-based PF analysis refers to developing NNs to approximate the state variables of power systems based on given historical system operation data, hereafter called dataset. The dataset includes input features $\vec{x}$ and output labels $\vec{y}$. For a power system with a reference bus and load buses, the input features are known variables, i.e., the net active and reactive power injections at load buses $\vec{x}=\{(\vec{p}^{d}_{i},\vec{q}^{d}_{i}):i=1,2,\dots,n\}$. The output labels are unknown variables, i.e., the real and imaginary components of complex voltages at load buses $\vec{y}=\{(\vec{\mu}_{i},\vec{\omega}_{i}):i=1,2,\dots,n\}$. The dimension of $\vec{x}$ and $\vec{y}$ is therefore $n\times 2$, where $n$ is the number of load buses. Note that the voltage magnitude and phase angle are known for the reference bus $i=0$, and the net active and reactive power injections are unknown. Having voltages at all load buses approximated, the net active and reactive power injections at the reference bus can be calculated.
The training of NNs is an iterative process and involves four steps. In the first iteration, the set of trainable parameters of the NN, i.e., the weight matrices and bias vectors $\theta=\{(W_{k},b_{k}):k=1,2,\dots,m\}$, are initialized in the update step, where $m$ is the number of hidden layers. In the forward step, the network is developed using (3b) and (4). The deviations of the approximated output obtained by the NN $\hat{\vec{y}}$ from the output labels $\vec{y}$ are computed in the loss step. Finally, the gradient of the deviations is calculated with respect to $\theta$ in the backward step. For the next iteration, $\theta$ is fine-tuned in the update step to reduce the deviations. The process continues until the maximum number of epochs is reached. In practice, the four steps can be individually and jointly modified to improve the overall performance of NNs. For example, it can be done by enhancing the network architecture in the forward step or adding a physical penalty term to the loss function in the loss step.